# LaVin-DiT: Large Vision Diffusion Transformer

### v3

@@ Abstract
This paper presents the Large Vision Diffusion Transformer (LaVin-DiT), a scalable and unified foundation model designed to tackle over 20 computer vision tasks in a generative framework. Unlike existing large vision models directly adapted from natural language processing architectures, which rely on less efficient autoregressive techniques and disrupt spatial relationships essential for vision data, LaVin-DiT introduces key innovations to optimize generative performance for vision tasks. First, to address the high dimensionality of visual data, we incorporate a spatial-temporal variational autoencoder that encodes data into a continuous latent space. Second, for generative modeling, we develop a joint diffusion transformer that progressively produces vision outputs. Third, for unified multi-task training, in-context learning is implemented. Input-target pairs serve as task context, which guides the diffusion transformer to align outputs with specific tasks within the latent space. During inference, a task-specific context set and test data as queries allow LaVin-DiT to generalize across tasks without fine-tuning. Trained on extensive vision datasets, the model is scaled from 0.1B to 3.4B parameters, demonstrating substantial scalability and state-of-the-art performance across diverse vision tasks. This work introduces a novel pathway for large vision foundation models, underscoring the promising potential of diffusion transformers. The code and models will be open-sourced.
@@ 1 Introduction
Large language models (LLMs) like GPT [10] and LLaMA [58] have rapidly gained widespread attention and transformed the field, demonstrating the strong capability to handle a wide range of language tasks within a unified framework [1]. This breakthrough of integrating diverse language tasks into a single large model has sparked momentum to develop similar large models for computer vision. The potential to create large vision models (LVMs) capable of generalizing across multiple vision tasks represents a promising step toward a more versatile, scalable, and efficient approach to vision-based AI [8, 39, 63, 5].
@@
However, constructing LVMs presents greater complexity than LLMs due to the inherently diverse and high-dimensional nature of vision data, as well as the need to handle variations in scale, perspective, and lighting across tasks [48, 31, 70, 52]. To handle the problem, recent work [5] has developed a sequential modeling method that learns from purely vision data by representing images, videos, and annotations in a unified “visual sentence” format. This method enables the model to predict sequential vision tokens from a vast dataset, entirely independent of language-based inputs (see Figure 1(a)). Although this method has shown promising results in diverse vision tasks, it faces two primary challenges. Specifically, the first issue concerns the efficiency limitations inherent in autoregressive sequence modeling [55], as it demands token-by-token prediction, which is computationally intensive for high-dimensional vision data [49]. The second issue is the disruption of spatial coherence when converting vision data into a sequential format, which compromises the preservation of spatial dependencies crucial for performance in vision tasks [77].
@@
In this paper, we introduce a large vision diffusion transformer (LaVin-DiT) to advance the development of next-generation LVMs. LaVin-DiT enjoys better computational efficiency and effectively preserves spatial relationships within vision data, thereby achieving superior performance across diverse vision tasks (see Figure 1(b)). Technically, to tackle the high-dimensional nature of vision data, we introduce a spatial-temporal variational autoencoder [30] that encodes data (i.e., image and video) into a continuous latent space, allowing compact representation while preserving essential spatial and temporal features. This reduces computational demands and improves efficiency without sacrificing the model’s ability to capture complex patterns. Besides, for generative modeling, we augment an existing diffusion transformer and propose a joint diffusion transformer with full-sequence joint attention. This module synthesizes visual outputs through parallel denoising steps, effectively reducing sequential dependencies to enhance processing efficiency while maintaining the spatial coherence essential for vision tasks. Moreover, to support unified multi-task training [61], we incorporate in-context learning [10, 19, 65, 73], where input-target pairs guide the diffusion transformer in aligning outputs with specific tasks. During inference, LaVin-DiT leverages task-specific context sets and test data as queries to adapt to various tasks without fine-tuning. This capability enables LaVin-DiT to achieve robust generalization across diverse tasks, leading to a versatile solution for complex vision applications.
@@ 2 Related Work
The image-resembling generation methods reformulate visual tasks as image generation problems, which allows models to handle dense visual predictions through inpainting and reconstruction tasks [8]. For instance, Painter [63] formulates dense prediction tasks as masked image inpainting, demonstrating in-context capabilities across multiple vision tasks. By leveraging pre-trained diffusion models [49], several methods [64, 24, 43, 60] utilize visual or textual instruction to guide generation and enhance adaptability across various tasks. The sequential modeling methods are largely inspired by breakthroughs in large language models and apply the sequence-to-sequence framework to visual data [58]. For these methods, visual data is typically quantized into sequences of discrete tokens [59]. The model is optimized through next-token prediction [10]. Recently, Bai et al. [5] introduce a framework that extends this concept to vision without relying on linguistic data, which treats visual data as a “visual sentence”. By representing images and videos as one-dimensional sequences, this method [5] enables a unified transformer that can tackle image and video tasks within a single framework, expanding the scope of sequential modeling in computer vision.

### v4

@@ Abstract
This paper presents the Large Vision Diffusion Transformer (LaVin-DiT), a scalable and unified foundation model designed to tackle over 20 computer vision tasks in a generative framework. Unlike existing large vision models directly adapted from natural language processing architectures, which rely on less efficient autoregressive techniques and disrupt spatial relationships essential for vision data, LaVin-DiT introduces key innovations to optimize generative performance for vision tasks. First, to address the high dimensionality of visual data, we incorporate a spatial-temporal variational autoencoder that encodes data into a continuous latent space. Second, for generative modeling, we develop a joint diffusion transformer that progressively produces vision outputs. Third, for unified multi-task training, in-context learning is implemented. Input-target pairs serve as task context, which guides the diffusion transformer to align outputs with specific tasks within the latent space. During inference, a task-specific context set and test data as queries allow LaVin-DiT to generalize across tasks without fine-tuning. Trained on extensive vision datasets, the model is scaled from 0.1B to 3.4B parameters, demonstrating substantial scalability and state-of-the-art performance across diverse vision tasks. This work introduces a novel pathway for large vision foundation models, underscoring the promising potential of diffusion transformers. The code and models are available.
@@ 1 Introduction
Large language models (LLMs) like GPT [10] and LLaMA [60] have rapidly gained widespread attention and transformed the field, demonstrating the strong capability to handle a wide range of language tasks within a unified framework [1]. This breakthrough of integrating diverse language tasks into a single large model has sparked momentum to develop similar large models for computer vision. The potential to create large vision models (LVMs) capable of generalizing across multiple vision tasks represents a promising step toward a more versatile, scalable, and efficient approach to vision-based AI [8, 40, 65, 5, 41].
@@
However, constructing LVMs presents greater complexity than LLMs due to the inherently diverse and high-dimensional nature of vision data, as well as the need to handle variations in scale, perspective, and lighting across tasks [50, 31, 75, 54, 68, 66]. To handle the problem, recent work [5] has developed a sequential modeling method that learns from purely vision data by representing images, videos, and annotations in a unified “visual sentence” format. This method enables the model to predict sequential vision tokens from a vast dataset, entirely independent of language-based inputs (see Figure 1(a)). Although this method has shown promising results in diverse vision tasks, it faces two primary challenges. Specifically, the first issue concerns the efficiency limitations inherent in autoregressive sequence modeling [57], as it demands token-by-token prediction, which is computationally intensive for high-dimensional vision data [51, 70]. The second issue is the disruption of spatial coherence when converting vision data into a sequential format, which compromises the preservation of spatial dependencies crucial for performance in vision tasks [83].
@@
In this paper, we introduce a large vision diffusion transformer (LaVin-DiT) to advance the development of next-generation LVMs. LaVin-DiT enjoys better computational efficiency and effectively preserves spatial relationships within vision data, thereby achieving superior performance across diverse vision tasks (see Figure 1(b)). Technically, to tackle the high-dimensional nature of vision data, we introduce a spatial-temporal variational autoencoder [30] that encodes data (i.e., image and video) into a continuous latent space, allowing compact representation while preserving essential spatial and temporal features. This reduces computational demands and improves efficiency without sacrificing the model’s ability to capture complex patterns. Besides, for generative modeling, we augment an existing diffusion transformer and propose a joint diffusion transformer with full-sequence joint attention. This module synthesizes visual outputs through parallel denoising steps, effectively reducing sequential dependencies to enhance processing efficiency while maintaining the spatial coherence essential for vision tasks. Moreover, to support unified multi-task training [63], we incorporate in-context learning [10, 19, 69, 78, 34], where input-target pairs guide the diffusion transformer in aligning outputs with specific tasks. During inference, LaVin-DiT leverages task-specific context sets and test data as queries to adapt to various tasks without fine-tuning. This capability enables LaVin-DiT to achieve robust generalization across diverse tasks, leading to a versatile solution for complex vision applications.
@@ 2 Related Work
The image-resembling generation methods reformulate visual tasks as image generation problems, which allows models to handle dense visual predictions through inpainting and reconstruction tasks [8]. For instance, Painter [65] formulates dense prediction tasks as masked image inpainting, demonstrating in-context capabilities across multiple vision tasks. By leveraging pre-trained diffusion models [51, 79], several methods [67, 24, 45, 62] utilize visual or textual instruction to guide generation and enhance adaptability across various tasks. The sequential modeling methods are largely inspired by breakthroughs in large language models and apply the sequence-to-sequence framework to visual data [60]. For these methods, visual data is typically quantized into sequences of discrete tokens [61]. The model is optimized through next-token prediction [10]. Recently, Bai et al. [5] introduce a framework that extends this concept to vision without relying on linguistic data, which treats visual data as a “visual sentence”. By representing images and videos as one-dimensional sequences, this method [5] enables a unified transformer that can tackle image and video tasks within a single framework, expanding the scope of sequential modeling in computer vision.